# Multi-Level Embedding and Alignment Network with Consistency and Invariance Learning for Cross-View Geo-Localization

### v2

@@ Abstract
Cross-View Geo-Localization (CVGL) involves determining the localization of aerial images by retrieving the most similar GPS-tagged satellite images. However, the imaging gaps between platforms are often significant and the variations in viewpoints are substantial, which limits the ability of existing methods to effectively associate cross-view features and extract consistent and invariant characteristics. Moreover, existing methods often overlook the problem of increased computational and storage requirements when improving model performance. To handle these limitations, we propose a lightweight enhanced alignment network, called the Multi-Level Embedding and Alignment Network (MEAN). The MEAN framework uses a progressive multi-level enhancement strategy, global-to-local associations, and cross-domain alignment, enabling feature communication across levels. This allows MEAN to effectively connect features at different levels and learn robust cross-view consistent mappings and modality-invariant features. Moreover, MEAN adopts a shallow backbone network combined with a lightweight branch design, effectively reducing parameter count and computational complexity. Experimental results on the University-1652 and SUES-200 datasets demonstrate that MEAN reduces parameter count by 62.17% and computational complexity by 70.99% compared to state-of-the-art models, while maintaining competitive or even superior performance. Our code and models will be released on https://github.com/ISChenawei/MEAN.
@@
Cross-view geo-location, invariance and consistency, progressive multi-level augmentation, cross-domain alignmentI Introduction
Cross-view geo-localization (CVGL) has received extensive attention in autonomous vehicles, aerial photography, and autonomous navigation [1, 2, 3]. CVGL is usually recognized as an image retrieval task on the heterogeneous platform to accurately determine the geo-location of a query image by matching it with several reference images taken from the varying views of different platforms. The early applications were concentrated on the matching of ground panoramic images with satellite images [4, 5, 6]. Recently, drone aerial photography has further extended its applications including drone target location and drone navigation [7, 8, 9]. In these cases, the aerial images taken by drones can be matched with satellite images of the same geographic area with precise longitude and latitude coordinates, achieving indirect drone location. This realizes drone navigation and replaces the need for global navigation satellite system (GNSS) equipment onboard [10]. However, CVGL encounters substantial challenges due to appearance variations and spatial misalignments caused by imaging gaps, scale variations and spatial transformations in the aerial photography scene.
@@
Although these methods have made significant progress in addressing feature distribution inconsistencies through dense partition learning and feature alignment, they often overemphasized fine-grained feature alignment. In this case, it relied heavily on the specific details of the viewpoints or scales, making it difficult to effectively associate features across different levels. In scenarios characterized by substantial viewpoint differences or large-scale spatial variations, consistency mappings may not capture the critical cross-view invariant that differs from these environmentally sensitive features, i.e., the invariance of features [25]. This limited their robustness and uniformity, ultimately undermining the generalization capability.
To address those issues, we propose a novel enhanced alignment network for the cross-view geo-localization tasks, called multi-level embedding alignment network (MEAN). In the proposed MEAN, ConvNeXt-Tiny, the smallest network in the ConvNeXt family [26] is used as a backbone to extract coarse-grained features. These features are then processed separately by the three designed branches. The progressive diversification embedding branch proposed in the MEAN adopts a stepwise expansion strategy to generate a rich set of embedding features, allowing the model to learn diverse feature representations. Furthermore, it uses contrastive loss to enhance the consistency and discriminability of the shared embeddings. To better coordinate feature representation capabilities, the other branch, named the global expansion embedding branch, jointly optimizes global and fine-grained feature representations to realize the global-to-local associations. Additionally, the multi-level feature fusion and adaptive calibration strategies with multi-level constraint achieve precise and robust feature alignment in the embedding space in the cross-domain enhanced alignment branch. Therefore, the proposed MEAN not only enhances feature representation, but also deeply explores the invariant modes and potential commonalities within the features. As shown in Fig. 1, our method significantly achieves improvements in both matching accuracy and computational efficiency compared to existing state-of-the-art methods. Its main contributions are summarized as follows.
@@ II RELATED WORKS
In this section, we provide a concise overview of two main methods commonly employed in cross-view geo-localization tasks: 1) feature extraction and contextual enhancement, and 2) feature alignment and optimization for discriminative power.
II-A Feature Extraction and Context Enhancement
@@
The first category of methods focused primarily on enhancing feature representation, typically achieving cross-view geo-localization through single global or local feature extraction. LCM [11] mapped aerial and satellite images into a unified feature space, transforming the task into a global location classification to learn the overall structure of the image. However, the ability to learn fine-grained features was limited in such global methods. Therefore, LPN [13] introduced a local pattern partitioning method that segmented images into multiple distance-based regions using a square ring partitioning strategy. This method effectively extracted fine-grained contextual information surrounding the target. Although this local method improved the model’s sensitivity to detail, it was found to be limited in handling significant viewpoint and scale variations by the extraction of the single global or local feature. Moreover, relying on a single branch for processing either global or local features hindered the effective integration of multi-scale and multi-level spatial information in cross-view scenarios. To address these issues, IFSs [8] proposed a multi-branch joint representation learning strategy that integrated global and local feature information through a multi-branch structure.
@@
However, these methods were mainly based on convolutional neural networks (CNNs), which are inherently limited in modeling long-range dependencies and complex contextual semantics, particularly in cross-view geo-localization tasks [22]. Transformer and ConvNeXt architectures have been gradually adopted in cross-view geo-localization due to their advantages in long-range dependency modeling and contextual information extraction. TransFG [10] used Transformer-based feature aggregation and gradient-guided modules to effectively integrate global and local information. SRLN [27] based on Swin Transformer [28] combined direction guidance and multi-scale feature fusion strategies, effectively bridging viewpoint and scale discrepancies. MCCG [7], on the contrary, introduced a multi-classifier method based on ConvNeXt to learn rich feature representations. However, the emphasis on contextual information can lead to over-attention to non-essential features. To mitigate this, CCR [29] introduced counterfactual causal reasoning to strengthen the model’s attention mechanism, which is useful to distinguish between essential and non-essential features. MFJR [17] used a multi-level feedback joint representation learning method, incorporating an adaptive region elimination strategy to effectively suppress irrelevant information and focus on key target features.
@@
II-B Feature Alignment and Discriminative Optimization
To achieve consistency in extracting features across different cross-views, researchers have utilized spatial alignment strategies to learn more consistent and discriminative features within a shared feature space. PCL [23] utilized Perspective Projection Transformation (PPT) to align aerial and satellite images to reduce spatial misalignment between views, and then employed a CGAN to synthesize realistic satellite image styles to narrow the imaging gap. This explicit alignment method struggled to learn internal feature differences effectively and might add unnecessary noise. Consequently, many researchers have attempted to integrate the metric loss and the contrastive loss within a shared space to identify the internal discrepancies. However, this method is often challenged by interference from viewpoint variations and exhibits limited discriminative power for negative samples. To address this, Sample4Geo [5] employed a hard negative sampling strategy to improve the model’s feature discrimination and contrastive learning effectiveness. However, significant differences in feature representations across cross-views posed challenges to the extraction and alignment of geographical features under substantial viewpoint variations. Relying solely on constraint-based alignment within a shared embedding space has been shown to be insufficient to address these discrepancies. Therefore, CAMP [30] introduced contrastive attribute mining and position-aware partitioning strategies to align geographic features under varying viewpoints and scales. Although this method has enhanced local feature consistency and discrimination, it excelled in extracting explicit differences across viewpoints but struggled to learn feature invariance under significant changes in viewpoint and scale, thereby limiting the model’s generalization capability. DAC [19] adopted domain alignment and scene consistency constraints to achieve coarse-to-fine feature consistency, relying primarily on direct alignment without fully exploring deep invariant patterns across viewpoints. As a consequence, DAC encountered challenges in maintaining stable feature mappings under extreme variations in viewpoint or scale. Furthermore, both CAMP and DAC incurred high computational costs in enhancing alignment accuracy, thereby affecting model efficiency and limiting deployment flexibility.
@@ III Multi-Level Embedding and Alignment Network
Problem Formulation: Given a cross-view geo-localization dataset, we denote the input image pairs as , where and represent the images taken from drone and satellite views, respectively. Each image is associated with a label , where denotes the total number of categories corresponding to different locations in the dataset. For example, in the University-1652 dataset, there are 701 buildings, each containing one satellite view image and multiple drone view images. For the cross-view geo-localization task, the goal is to learn a mapping function that projects images from different platforms (e.g., drone and satellite) into a shared semantic space. In this space, the features of images representing the same location should be close to each other, while those from different locations should be well separated. This method enables effective matching of images from different views in the same location by leveraging feature similarity. It can address the challenges of significant viewpoint differences and supports accurate geo-localization.
@@ III-A ConvNeXt-Tiny Backbone for Feature Extraction
In this work, we use ConvNeXt-Tiny as the backbone for the extraction of features. ConvNeXt-Tiny is the most lightweight variant within the ConvNeXt family, combining depthwise convolutions and layer normalization to provide an efficient CNN solution for cross-view geo-localization tasks.
ConvNeXt-Tiny processes input images with a resolution of 384×384 pixels, and the output feature maps are represented as follows,
@@
Compared with deeper and more complex backbone networks such as Transformer-based [17] and ConvNeXt-based [29] models for cross-view geo-localization, ConvNeXt-Tiny reduces the parameter count by 67%. However, because of the shallow structure and compact design, it has certain limitations to learning complex and deep information about features. To address these shortcomings, we introduce complementary branches, designed as follows.
@@ III-B Progressive Extension Embedding branch
DEG module: DEG extracts contextual information from multiple embeddings through parallel pathways, enhancing representational consistency in different embeddings. As illustrated in Fig.2, for the output of the feature map by the backbone network , three dilated convolution layers with a kernel size of are applied. These layers, namely , , and have dilation rates of 1, 2 and 3, respectively. They reduce the channel size of the feature map to one-fourth of its original size. The feature maps are then fused by combining the outputs into a single feature map, followed by the ReLU activation layer to improve the nonlinearity of the DEG’s representation. Subsequently, another convolutional layer with a kernel size of , denoted as , is applied to the resulting feature map to restore its dimension to match that of the original feature map . Based on this, we obtain compact and enhanced embedded features with improved diversity . To adaptively address feature alignment and enhancement across different task scenarios, we introduce a feature augmentation and balanced fusion strategy. This strategy combines the generated embedding features with the original features through weighted fusion, preserving essential information from the original features while enhancing the overall feature representation. Finally, the DEG module generates two identical feature representations, both of which undergo a dropout operation to mitigate overfitting and enhance the model generalization capability. These two sets of feature outputs will be used in subsequent classification procedure to expand the discriminative power of the model and enable it to co-optimize across different feature subspaces. The feature embedding is represented as follows,
@@
Subsequently, adaptive average pooling (AvgPool) is applied to aggregate the multi-scale information, producing more discriminative semantic embeddings. Finally, the compact feature representation is processed through the batch normalization operation , followed by a dropout operation for regularization to aviod overfitting, and then propagated to the linear layer for classification. We outline these processing operations , , , as a lump
@@ III-C Global Extension Embedding Branch
The GEE branch adopts a global context aggregation strategy compared to the PEE branch, which progressively generates and enhances local embedded details. In the GEE branch, a mean pooling operation replaces the DEG module, processing the output features of the backbone to obtain a global feature representation. The global feature representation is then propagated to the DEC module for further refinement.
By applying the mean pooling to , the GEE branch provides a more global perspective of feature representation before diverse embedded local detail processing. The backbone output feature undergoes mean pooling to compute the global average feature representation as follows,
where, and are height factor and width factor, respectively. This operation compresses the spatial dimensions, where . The global feature is then fed into the DEC module for further processing to obtain described as the DEC module III-B. In this way, the GEE branch can aggregate global information, generating comprehensive global feature embeddings that provide the DEC module with context-aware global features. Based on this global perspective, global and local consistency is ensured in the GEE branch.

### v4

@@ Abstract
Cross-View Geo-Localization (CVGL) involves determining the localization of aerial images by retrieving the most similar GPS-tagged satellite images. However, the imaging gaps between platforms are often significant and the variations in viewpoints are substantial, which limits the ability of existing methods to effectively associate cross-view features and extract consistent and invariant characteristics. Moreover, existing methods often overlook the problem of increased computational and storage requirements when improving model performance. To handle these limitations, we propose a lightweight enhanced alignment network, called the multi-level embedding and alignment network (MEAN). The MEAN framework uses a progressive multi-level enhancement strategy, global-to-local associations, and cross-domain alignment, enabling feature communication across levels. This allows MEAN to effectively connect features at different levels and learn robust cross-view consistent mappings and cross-view invariant features. Moreover, MEAN adopts a shallow backbone network combined with a lightweight branch design, effectively reducing parameter count and computational complexity. Experimental results on the University-1652 and SUES-200 datasets demonstrate that MEAN reduces parameter count by 62.17% and computational complexity by 70.99% compared with state-of-the-art models, while maintaining competitive or even superior performance. Our code is available at https://github.com/ISChenawei/MEAN
@@
Cross-view geo-localization, Invariance and consistency, Progressive Multi-level augmentation, Cross-domain alignmentI Introduction
Cross-view geo-localization (CVGL) has received extensive attention in autonomous vehicles, aerial photography, and autonomous navigation [1, 2, 3]. CVGL is usually recognized as an image retrieval task on the heterogeneous platform to accurately determine the geo-location of a query image by matching it with several reference images captured from the varying views of different platforms. The early applications were concentrated on the matching of ground panoramic images with satellite images [4, 5, 6]. In recent years, with the continuous advancement of remote sensing technology [7, 8, 9], the applications of drone aerial photography have been further expanded, encompassing drone target localization, drone navigation, and other related fields [10, 11, 12]. In these cases, the aerial images captured by drones can be matched with satellite images of the same geographic area with precise longitude and latitude coordinates, achieving indirect drone location. This realizes drone navigation and replaces the need for global navigation satellite system (GNSS) equipment onboard [13]. However, CVGL encounters substantial challenges due to appearance variations and spatial misalignment caused by imaging gaps, scale variations and spatial transformations in the aerial photography scene.
@@
Although these methods have made significant progress in addressing feature distribution inconsistencies through dense partition learning and feature alignment, they often overemphasized fine-grained feature alignment. In this case, it relied heavily on the details of the viewpoints or scales, making it difficult to effectively associate features across different levels. In scenarios characterized by substantial viewpoint differences or large-scale spatial variations, consistency mappings may not capture the critical cross-view invariant that differs from these environmentally sensitive features, i.e., the invariance of features [28]. This limited their robustness and uniformity, ultimately undermining the generalization capability.
To address those issues, we propose a novel enhanced alignment network for the CVGL tasks, called multi-level embedding alignment network (MEAN). In the proposed MEAN, ConvNeXt-Tiny, the smallest network in the ConvNeXt family [29] is used as a backbone to extract coarse-grained features. These features are then processed separately by the three designed branches. The progressive diversification embedding branch proposed in the MEAN adopts a stepwise expansion strategy to generate a rich set of embedding features, allowing the model to learn diverse feature representations. Furthermore, it uses contrastive loss to enhance the consistency and discriminability of the shared embeddings. To better coordinate feature representation capabilities, the other branch, named the global expansion embedding branch, jointly optimizes global and fine-grained feature representations to realize the global-to-local associations. Additionally, the multi-level feature fusion and adaptive calibration strategies with multi-level constraint achieve precise and robust feature alignment in the embedding space in the cross-domain enhanced alignment branch. Therefore, the proposed MEAN not only enhances feature representation, but also deeply explores the invariant modes and potential commonalities within the features. As shown in Fig. 1, our method significantly achieves improvements in both matching accuracy and computational efficiency compared with existing state-of-the-art methods. Its main contributions are summarized as follows:
@@
The remainder of this paper is organized as follows. Section II systematically reviews prior research. In Section III, the proposed MEAN is presented in detail. The experimental results are reported and analyzed in Section IV. In Section V, a comprehensive analysis of the proposed MEAN is provided. Finally, conclusions are outlined in Section VI.
@@ II RELATED WORKS
In this section, we provide a concise review of the related work on CVGL, with an emphasis on two main methods commonly employed in CVGL tasks: 1) feature extraction and contextual enhancement, and 2) feature alignment and optimization for discriminative power.
II-A Cross-View Geo-Localization
CVGL has been generally regarded as an image retrieval task involving heterogeneous viewpoints, primarily focusing on ground-to-satellite and drone-to-satellite image retrieval. Early studies relied on handcrafted feature operators to extract and align features across distinct viewpoints [30, 31]. With the rapid advancement of deep convolutional neural networks (CNNs), recent works [32, 33, 34, 35] have shifted toward deep representation learning to learn cross-view features. This progress brought creation of multiple public datasets featuring ground–satellite image pairs such as CVUSA [36], CVACT [37]. A pre-trained CNN for high-level feature extraction in CVGL was first employed to demonstrate that the features encoded the semantic cues related to geographic location [38]. Subsequently, NetVLAD [39] was incorporated into a Siamese-like architecture to achieve robust image descriptors with significant viewpoint variation [36].
CVGL has been further extended to the drone-view, along with several specialized drone-view localization datasets. The University-1652 dataset has contained drone–satellite image pairs and reformulated the retrieval problem within a classification framework [1]. Subsequent studies expanded on this dataset by incorporating diverse environmental conditions, such as fog, rain, and snow, to systematically assess model robustness under real-world scenarios [40]. Additionally, the SUES-200 dataset [41] was proposed to specifically examine the influence of varying flight altitudes on CVGL performance.
II-B Feature Extraction and Context Enhancement
@@
The first category of methods focused primarily on enhancing feature representation, typically achieving CVGL through single global or local feature extraction. LCM [14] mapped aerial and satellite images into a unified feature space, transforming the task into a global location classification to learn the overall structure of the image. However, the ability to learn fine-grained features was limited in such global methods. Therefore, LPN [16] introduced a local pattern partitioning method that segmented images into multiple distance-based regions using a square ring partitioning strategy. This method effectively extracted fine-grained contextual information surrounding the target. Although this local method improved the model’s sensitivity to detail, it was found to be limited in handling significant viewpoint and scale variations by the extraction of the single global or local feature. Moreover, relying on a single branch for processing either global or local features hindered the effective integration of multi-scale and multi-level spatial information in cross-view scenarios. To address these issues, IFSs [11] proposed a multi-branch joint representation learning strategy that integrated global and local feature information through a multi-branch structure.
@@
However, these methods were mainly based on convolutional neural networks (CNNs), which are inherently limited in modeling long-range dependencies and complex contextual semantics, particularly in CVGL tasks [25]. Transformer and ConvNeXt architectures have been gradually adopted in CVGL due to their advantages in long-range dependency modeling and contextual information extraction. TransFG [13] used Transformer-based feature aggregation and gradient-guided modules to effectively integrate global and local information. SRLN [42] based on Swin Transformer [43] combined direction guidance and multi-scale feature fusion strategies, effectively bridging viewpoint and scale discrepancies. MCCG [10], on the contrary, introduced a multi-classifier method based on ConvNeXt to learn rich feature representations. However, the emphasis on contextual information can lead to over-attention to non-essential features. To mitigate this, CCR [44] introduced counterfactual causal reasoning to strengthen the model’s attention mechanism, which is useful to distinguish between essential and non-essential features. MFJR [20] used a multi-level feedback joint representation learning method, incorporating an adaptive region elimination strategy to effectively suppress irrelevant information and focus on key target features.
@@
II-C Feature Alignment and Discriminative Optimization
To achieve consistency in extracting features across different cross-views, researchers have utilized spatial alignment strategies to learn more consistent and discriminative features within a shared feature space. PCL [26] utilized perspective projection transformation to align aerial and satellite images to reduce spatial misalignment between views, and then employed a CGAN to synthesize realistic satellite image styles to narrow the imaging gap. This explicit alignment method struggled to learn internal feature differences effectively and might add unnecessary noise. Consequently, many researchers have attempted to integrate the metric loss and the contrastive loss within a shared space to identify the internal discrepancies. However, this method is often challenged by interference from viewpoint variations and exhibits limited discriminative power for negative samples. To address this, Sample4Geo [5] employed a hard negative sampling strategy to improve the model’s feature discrimination and contrastive learning effectiveness. However, significant differences in feature representations across cross-views posed challenges to the extraction and alignment of geographical features under substantial viewpoint variations. Relying solely on constraint-based alignment within a shared embedding space has been shown to be insufficient to address these discrepancies. Therefore, CAMP [45] introduced contrastive attribute mining and position-aware partitioning strategies to align geographic features under varying viewpoints and scales. Although this method has enhanced local feature consistency and discrimination, it excelled in extracting explicit differences across viewpoints but struggled to learn feature invariance under significant changes in viewpoint and scale, thereby limiting the model’s generalization capability. DAC [22] adopted domain alignment and scene consistency constraints to achieve coarse-to-fine feature consistency, relying primarily on direct alignment without fully exploring deep invariant patterns across viewpoints. As a consequence, DAC encountered challenges in maintaining stable feature mappings under extreme variations in viewpoint or scale. Furthermore, both CAMP and DAC incurred high computational costs in enhancing alignment accuracy, thereby affecting model efficiency and limiting deployment flexibility.
@@ III Multi-Level Embedding and Alignment Network
Problem Formulation: Given a CVGL dataset, we denote the input image pairs as , where and represent the images captured from drone and satellite views, respectively. Each image is associated with a label , where denotes the total number of categories corresponding to different locations in the dataset. For example, in the University-1652 dataset, there are 701 buildings, each containing one satellite view image and multiple drone view images. For the CVGL task, the goal is to learn a mapping function that projects images from different platforms (e.g., drone and satellite) into a shared semantic space. In this space, the features of images representing the same location should be close to each other, while those from different locations should be well separated. This method enables effective matching of images from different views in the same location by leveraging feature similarity. It can address the challenges of significant viewpoint differences and supports accurate geo-localization.
@@ III-A ConvNeXt-Tiny Backbone for Feature Extraction
In this work, we use ConvNeXt-Tiny as the backbone for the extraction of features. ConvNeXt-Tiny is the most lightweight variant within the ConvNeXt family, combining depthwise convolutions and layer normalization to provide an efficient CNN solution for CVGL tasks.
ConvNeXt-Tiny processes input images with a resolution of 384×384 pixels, and the output feature maps are represented as follows:
@@
Compared with deeper and more complex backbone networks such as Transformer-based [20] and ConvNeXt-based [44] models for CVGL, ConvNeXt-Tiny reduces the parameter count by 67%. However, because of the shallow structure and compact design, it has certain limitations to learning complex and deep information about features. To address these shortcomings, we introduce complementary branches, designed as follows.
@@ III-B Progressive Extension Embedding branch
DEG module: DEG extracts contextual information from multiple embeddings through parallel pathways, enhancing representational consistency in different embeddings. As illustrated in Fig.2, for the output of the feature map by the backbone network , three dilated convolution layers with a kernel size of are applied. These layers, namely , , and have dilation rates of 1, 2 and 3, respectively. They reduce the channel size of the feature map to one-fourth of its original size. The feature maps are then fused by combining the outputs into a single feature map, followed by the ReLU activation layer to improve the nonlinearity of the DEG’s representation. Subsequently, another convolutional layer with a kernel size of , denoted as , is applied to the resulting feature map to restore its dimension to match that of the original feature map . Based on this, we obtain compact and enhanced embedded features with improved diversity . To adaptively address feature alignment and enhancement across different task scenarios, we introduce a feature augmentation and balanced fusion strategy. This strategy combines the generated embedding features with the original features through weighted fusion, preserving essential information from the original features while enhancing the overall feature representation. Finally, the DEG module generates two identical feature representations, both of which undergo a dropout operation to mitigate overfitting and enhance the model generalization capability. These two sets of feature outputs will be used in subsequent classification procedure to expand the discriminative power of the model and enable it to co-optimize across different feature subspaces. The feature embedding is represented as follows:
@@
Subsequently, adaptive average pooling (AvgPool) is applied to aggregate the multi-scale information, producing more discriminative semantic embeddings. Finally, the compact feature representation is processed through the batch normalization operation , followed by a dropout operation for regularization to avoid overfitting, and then propagated to the linear layer for classification. We outline these processing operations , , , as a lump :
@@ III-C Global Extension Embedding Branch
The GEE branch adopts a global context aggregation strategy compared with the PEE branch, which progressively generates and enhances local embedded details. In the GEE branch, a mean pooling operation replaces the DEG module, processing the output features of the backbone to obtain a global feature representation. The global feature representation is then propagated to the DEC module for further refinement.
By applying the mean pooling to , the GEE branch provides a more global perspective of feature representation before diverse embedded local detail processing. The backbone output feature undergoes mean pooling to compute the global average feature representation as follows:
where and are height factor and width factor, respectively. This operation compresses the spatial dimensions, where . The global feature is then fed into the DEC module for further processing to obtain described as the DEC module III-B. In this way, the GEE branch can aggregate global information, generating comprehensive global feature embeddings that provide the DEC module with context-aware global features. Based on this global perspective, global and local consistency is ensured in the GEE branch.